# https://cs.uwaterloo.ca/~smwatt

### v1

@@ 4 Issues Encountered in Lisp Evolution
What fun. So let me feed in the following from my local position: (1) The first Lisp I (+ John Fitch) implemented so we could run Reduce was “Cambridge Lisp” and was coded in BCPL. The data representation used high bits in a word to tag things (and the IBM mainframe in use used 24-bit addressing so the top 8 bits were free..). For that our bignums used $10^{9}$ as a base because we argued that doing that made only a linear change in the cost of arithmetic but made reading and printing linear not quadratic. Numbers were never going to be big enough to mess with $n\log n$ schemes!
@@
After a significant while it became clear that BCPL was becoming less central to the world and that byte- not word-addressed machines were very dominant. So CSL (Standard Lisp coded in C) happened. By then C was ANSI C so it was not as nasty as K&R had been. That made bignums out of 31-bit chunks because C did not give trivial access to the carry bit. Data was tagged in the low 3 bits and complex data had a header word with more discrimination as well as a size field.
@@
Again the years passed. 64-bit machines arrived so I had to review all the code for 32-bit sensitivity. Bignums remained at 31-bit digits. The thing I think I will claim as most amazing was the checkpointing scheme. I checkpointed by writing a simple binary dump of the heap image and a word giving its start address. On reloading I parsed the heap well enough to relocate and for 64-bit use I could reload and map either way between 32 or 64-bit images. You might imagine that this was exploiting C with a detailed belief about storage representation (I could also allow for byte-sex [bit order], so a heap image made anywhere could be used anywhere else).
@@
Yet more years and I concluded that if eventually anybody was going to maintain this after me it needed to be in C++ not C. So I started by just renaming all the files. The C++ is much more fussy about “const char*”so there was a deal of work there. And I could now use “//” comments officially rather than “/* .. */” ones that where all older C systems certified as legal (whatever gcc actually coped with).
Since then I have gradually moved to use more C++ facilities. But the 2 bigger changes are that I ripped out the allocator and garbage collector and created a conservative one. And I have a new arithmetic package that will use 64-bit digits for bignums — but despite quite a few years in gestation that is not the standard one in place yet.
@@
For Reduce I continue to try to explain to people that when it was young you just said “it is coded in Lisp” and they thought they understood. But a better way to explain it now is that it has its own language and the parse trees for that and the raw data that is worked on happen to look very much like Lisp 1.6 — but that is an internal abstraction layer and almost nobody extending or maintaining Reduce work in Lisp as a source language. So the “lisp implementation” used stands a good chance of really being the “Reduce kernel” and not being used for any Lisp activity elsewhere. And the main versions used are thus customised and tuned to support Reduce. Hearn and his collaborators also found it helpful to update and extend the specification of the Lisp Standard[37] that they would rely on, so the variants cited here agreed to that new contract. “CSL”[39] is coded in C++, while “PSL”[12] builds on the Utah Portable Lisp Compiler[13] and is coded in “syslisp” which is a Lisp that has functions like (wplus a 1) that increments the machine representation of the value of a, and so that and its friends let it implement bignums and garbage collection rather as if it was a parse tree for C where “int c = a+b” has turned into “(prog (c) (setq c (wplus a b))...”.
@@
I also like assert to that the Reduce language is the successful version of Lisp 2[9]. McCarthy thought in terms of S-notation and M-notation and Lisp 2 was to support the latter (but the project basically failed). Lisp 2 “had largely Lisp-like semantics and ALGOL 60-like syntax”. The Reduce “rlisp” notation is a really rather close relative of what the Lisp 2 people designed — but it was made to work, work portability and has survived in use for very many years…. and one can see Scratchpad’s boot as very much a variant on the same concept.
@@
Note that Reduce and C were being invented at about the same times, but I think it took longer for C to be distributed beyond Bell Labs than for Reduce to get into circulation.
@@ 5 Case Study: Reduce
The code that eventually became known as Reduce was originally coded directly in Lisp 1.5 — but by the end of the 1960s it had been reworked so that it was written in its own “Algol-like” syntax that for the system programmer offered Lisp-like semantics. Building the system called for a fairly modest size parser in regular Lisp to bootstrap things until the system’s better and more complete version of the same parser expressed in its own language could be installed. This scheme provided significant isolation from dialect variations between competing Lisp implementations — the bootstrap parser could need to be adapted to allow for differences in support for character handling, but then the transformation from textual source code to internal s-expressions readily provided a location where adjustments to the use of Lisp functions could be made. To give a concrete example of the sort of adjustment needed, we note that the Lisp operators map and mapcar that apply a given function to each component of a list took their arguments in differing orders basically depending on whether the Lisp implementation concerned followed East Coast or West Coast traditions. Reduce could adapt for that by including conversion code as part of its parsing phase.
@@
Over time and as the number of computers (and hence Lisp systems) that people wanted to run Reduce on grew this scheme which amounted to source to source translation from the dialect Reduce expected to whatever was locally available became clumsier and more onerous. In response to this Reduce codified a Lisp dialect it would use (calling it “Standard Lisp”[15] so that host systems could arrange to provide a compatibility layer that exposed this agreed version, and the code within Reduce being written based on that could then just run without need for further change. To make this process as easy as possible Standard Lisp mandated only facilities it expected that all plausible Lisps would be able to support, so it did not provide any of the more elaborate additions to Lisp that those who were using the language directly were developing. For instance more or less elaborate loop constructs were present in the Reduce source language but got expanded to simple forms by the time a Lisp “parse tree“ had been generated, so there was no need for the supporting Lisp to provide clever iteration functions.
@@
So the increasing bulk of Reduce being coded in its own language and sitting on a rather minimal base really helped with portability, but two challenges could arise. One was that newer parts of the system sometimes needed facilities not captured in Standard Lisp. Perhaps the best example to quote would be modular arithmetic. This could be provided by merely defining all the necessary functions in terms of ordinary arithmetic with a lot of calls to the remainder function, but really it would be better if implemented at a lower level. Well Reduce could cope with that! It could include portable reference implementations of everything is needed, but a particular implementation could offer built-in and presumably better versions and then tag the relevant function names as “lose” which would cause Reduce to skip instating its reference code. This strategy maintained the policy that the bulk of the code should only very very rarely need to be conditialised based on the platform it was running on.
@@
The second challenge was of course performance. Main-stream Lisps were increasingly focusing their implementation efforts on supporting the full richness of environment expected by users who were coding directly in the parenthesised notation and the optimizations they were delivering were not necessarily the ones most important to Reduce. So as a spin-off activity Reduce first developed its own explicitly portable Lisp compiler, and went on from there is develop of complete Lisp (PSL[12]) explicitly providing (a superset of) Standard Lisp, reasonably easy to move to new machines because it used the portable compiler, and particularly tuned to run Reduce well. PSL was essentially all coded in Lisp as extended with functions that were there to compile into simple operations on machine registers and memory. In broadly the same time-frame Cambridge Lisp[10] was implemented mostly in BCPL but also using a version of the Reduce project’s compiler, and BCPL’s focus on portability meant that ended up being used on a range of the emerging small computers from Amiga, Atari, Acorn and the like. In each case these distilled portability away from the parts of Reduce that supported algebra, and they could both focus on trying to be economical in memory and good in terms of speed for the particular load that Reduce placed on them.
@@
Somewhat later on the use of BCPL became less tenable as C took over the ecological niche it had inhabited and as Unix became a dominant operating system in academia. So Cambridge Lisp was replaced by CSL[39], which was a fully fresh Lisp this time coded in C. A particular feature of this was that its almost entire purpose was to act as a kernel for Reduce rather than as a Lisp for more general use. At the time memory often limited problem solving ability. To save space CSL compiled the Lisp that came out of the Reduce parser into a compact byte-coded form which it could then interpret. One initial insight was that the bytecode interpreter was liable to fit within the caches of the more important computers of the times, and that would lead to the overhead of interpretation being much less severe than one might have at first feared. The second performance trick was to profile a wide range of Reduce applications and on that basis identify the most crucial functions. Those were then compiled from Lisp into equivalent C code that followed the internal CSL conventions and got passed through the system C compiler and linked in as parts of the main CSL executable. By then the leading C compilers (notably gcc) were starting to provide optimization capabilities distinctly superior to the more generic ones present in the available Lisp compilers. The C code in CSL was kept fairly conservative with a view to keeping it as portable as was reasonably possible.
@@
Because this story spans many years, one should not be amazed to notice that just as BCPL had started to look long in the tooth, at some stage the C code in CSL started to be reworked into C++. The changes there forced further case regarding for instance the difference between immutable strings and updateable arrays of characters, the issues of “strict aliasing”, and the fact that arithmetic overflow on signed integers was deemed undefined. Responding to each of these will have tended to enhance some mix of reliability and portability of the code-base. One of the more concerning issues while that transition was going on was the proper delay time between C++ standards codifying useful new features and the moment that assuming their availability would not seriously compromise attempts to build the code on older machines. While it is often reasonable to expect that personal computers will be kept well up to date with the most recent versions of compilers and other software, major institutional systems can prioritise stability to an extent that to a developer who has not encountered that before it at first seems that they have hardly moved on from the ark. CSL has tried to steer clear of reliance on new language features until they have had plenty of time to propagate to all potential use-sites. This is again in line withe the long standing Reduce philosophy of preferring easy portability over extra and glossy facilities. Over its lifetime the storage allocation and garbage collection parts of CSL have been re-written several times, the scheme for preserving a system on disc for later re-loading has been changed also several times and the large integer arithmetic implementation altered rather substantially. But from the outside all these changes tried to be invisible so that the Reduce system sityting on top did not need to pay any attention but could benefit from speed improvements or reliability improvements.
@@ 6 Case Study: Aldor
Aldor purposefully retained most of the ideas and syntax of [29] while generalizing the type system and providing features to allow clear separation of base language and libraries. Functions and types were first-class values and the language design allowed efficient compilation. Dependent types provided genericity and the structures of modern algebra (such as rings, fields and modules) can be represented naturally as type categories. An optimizing compiler for Aldor could generate C code in the ANSI (C90) or K&R dialects, for stand-alone or linked programs, or Lisp in the Common Lisp, Standard Lisp or Scheme dialects, for embedding in systems based on those. For a fuller description, see [48, 49, 50]. Recall that during its development Axiom had been named Scratchpad II, and similarly Aldor was named $A^{\sharp}$ (and briefly as Axiom-XL).
The main development of Aldor spanned a period of about ten years, from 1984 to 1995. It began as an effort to produce an improved implementation of the language of [29]. It was written in the “boot” language, described earlier, and it compiled algebra code to Lisp to run within the Scratchpad II system.
@@
A wide range of operating systems were supported, including Microsoft DOS and Windows 3.1, IBM OS2 and CMS, Digital Equipment VAX VMS, Apple MAC System 7, and a wide range of Unix variants including AIX PS2, AIX RT, IX RS, AIX 370, AIX ESA (all from IBM), SunOS and Solaris from Sun Microsystems, the Open Software Foundation’s OSF/1, as well as Unicos, Irix, Next, Linux, Ultrix and HP UX. The C environments supported included those from Borland, Metaware, Metroworks, Microsoft, the Free Software Foundation (gcc), four from IBM, DEC’s VMS and Ultrix C, as well as those from Apple, Cray, Hewlett Packard, MIPS and Sun. The architectures supported included Intel 8086 and 80386, Motorolla 68000, PowerPC, IBM 370, RT and RS, DEC VAX and AXP, Sun Sparc, MIPS, Cray and Hewlet Packard PA.
@@
The portability layer provided a set of very low-cost abstractions to give a common coding interface for all the supported platforms. This allowed efficient low-level programming while shielding the main body of the code from defects in the various C compilers (of which there were many) and providing common concepts where operating systems differed. This approach necessitated a certain amount of system-specific implementation of low-level functions (about 5350 lines of C code for all the operating system variants listed above). The portability layer was the only layer in which platform-specific #ifdefs could occur.
@@
The portability layer also provided a range of macros to encapsulate code that might have different variants depending on the named properties. For example, ptrDiff(p,q) had different implementations depending on whether or not pointers were canonical. Strictly speaking, the C standard would not allow subtraction of pointers into different objects, so the ptrDiff operation was needed to support garbage collection. Another example would be that in some (but not all) environments, functions to be used as signal handlers had to have a special modifier in their declaration, and this modifier varied. So a macro SignalModifier would always be used in such declarations, and it would sometimes expand to a keyword and otherwise to nothing. All told, there were about 80 of these macros.
Although the C standard [21] was very new and support for it was lacking on many important platforms, the Aldor source code would be written to the standard. A software tool was written to convert function prototypes and new-style function headers if needed on a given platform, e.g. extern int f(char *s); and int f(char *s) { ... }, to K&R style, e.g. extern int f(); andint f(s) char *s; { ... }.
@@
Aldor’s run time library provided a conservative garbage collector that made use of the osMemMap function. The root set for garbage collection was obtained by scanning all data regions outside the Aldor heap for potential pointers into the Aldor heap.
@@
To give an idea of the amount of portable versus platform-specific code, we note the compiler and run-time system consisted of about 210,000 lines of C code, 1000 lines of 2-level grammar code, and 2,400 lines of Aldor code. Of the C code, about 8000 lines were devoted to the portability layer. About 2000 lines of C code were for software tools were used to build the compiler. These included oldc (to convert ANSI C to K&R C), zacc (to convert 2-level grammar to yacc), and atinlay (to convert the message database to pig latin, for multilingual testing). Libraries in the Aldor language completed the system, with about 11,000 lines providing basic types and functionality and an additional 41,000 lines providing symbolic mathematical functions.
@@
What we take from this experience is that, with due care and attention to detail, it has been possible to build a highly portable software system for computer algebra that runs across a wide range of diverse platforms by providing a low-level portability layer comprising about 3% of the total code base. The resulting system was able to produce code that ran at speeds comparable to hand-coded C or CC++ for the sorts of problems these systems could handle [48, 8].

### v4

@@ 4 Issues Encountered in Lisp Evolution
The first Lisp ACN and John Fitch implemented so they could run Reduce was “Cambridge Lisp” and was coded in BCPL. The data representation used high bits in a word to tag things (and the IBM mainframe in use used 24-bit addressing so the top 8 bits were free). For that the bignums used $10^{9}$ as a base because doing that made only a linear change in the cost of arithmetic but made reading and printing linear not quadratic. Numbers were never going to be big enough to mess with $n\log n$ schemes!
@@
After a significant while it became clear that BCPL was becoming less central to the world and that byte- not word-addressed machines were very dominant. So CSL (Standard Lisp was coded in C). By then C was Ansi C so it was not as nasty as K&R C had been. Bignums were made out of 31-bit chunks because C did not give trivial access to the carry bit. Data was tagged in the low 3 bits. Essentially small integers were represented as immediate values, cons-cells had a tag dedicated to mark references to them and all other data used a header word containing further discrimination as well as a size field.
@@
Again the years passed. 64-bit machines arrived so it was necessary to review all the code for 32-bit sensitivity. Bignums remained at 31-bit digits. The thing that was most amazing was the checkpointing scheme. A system image was checkpointed by writing a simple binary dump of the heap image and a word giving its start address. On reloading, the heap was parsed well enough to relocate and, for 64-bit use, it was possible to reload and map either way between 32 or 64-bit images. One might imagine that this was exploiting C with a detailed belief about storage representation (it also allowed for byte-sex [bit order], so a heap image made anywhere could be used anywhere else).
@@
Yet more years passed and ACN concluded that if eventually anybody else was going to maintain this software it needed to be in C++ not C. The C++ language is much more fussy about “const char*” so there was a deal of work there. It was now possible to use “//” comments officially rather than “/* .. */” ones that were all the older C systems certified as legal (whatever gcc actually coped with).
Since then, the implementation has gradually moved to use more C++ facilities. But some of the more radical changes may be worth noting:
  • The allocator and garbage collector were ripped out and replaced by a new conservative one. On 64-bit architectures, conservative collectors are less problematic — falsely identifying a data word as a pointer (and thus retaining garbage) is far more probable on 32-bit architectures.
  • A new arithmetic package has been written that will use 64-bit digits for bignums and which will be faster then the old one as well as cleaner code — but despite quite a few years in gestation that is not the standard one in place yet.
  • The checkpoint/restart scheme has been replaced with one using techniques akin to data serialization.
A point to be made about these and the many other changes that have been made at that level, is that they are isolated from the code that performs algebra as such and that they were motivated by the need to support Reduce on a broad range of modern systems and with the best possible performance and the best changes for future portability and maintenance.
@@
For Reduce one can explain to people that when it was young you just said “it is coded in Lisp” and they thought they understood. But a better way to explain it now is that it has its own language and the parse trees for that and the raw data that is worked on happen to look very much like Stanford Lisp 1.6 — but that it is an internal abstraction layer and that almost nobody extending or maintaining Reduce works in Lisp as a source language. So the “Lisp implementation” used stands a good chance of really being the “Reduce kernel” and not being used for any Lisp activity elsewhere. And the main versions used are thus customised and tuned to support Reduce. Hearn and his collaborators also found it helpful to update and extend the specification of the Lisp standard[38] that they would rely on, so the variants cited here agreed to that new contract. “CSL[40] is coded in C++, while “PSL[14] builds on the Utah Portable Lisp Compiler[15] and is coded in “syslisp” which is a Lisp that has functions like (wplus a 1) that increments the machine representation of the value of a, and so that and its friends let it implement bignums and garbage collection rather as if it was a parse tree for C where “int c = a+b” has turned into “(prog (c) (setq c (wplus a b))...
@@
It may also be asserted that the Reduce language is the successful version of Lisp 2[1]. McCarthy thought in terms of S-notation and M-notation and Lisp 2 was to support the latter (but the project basically failed). Lisp 2 “had largely Lisp-like semantics and ALGOL 60-like syntax”. The Reduce “rlisp” notation is a really rather close relative of what the Lisp 2 people designed — but it was made to work, work portably and has survived in use for very many years. One can see Scratchpad’s Boot as very much a variant on the same concept.
@@
Note that Reduce and C were being invented at about the same times, but it took longer for C to be distributed beyond Bell Labs than for Reduce to get into circulation.
@@ 5 Case Study: Reduce
The code that eventually became known as Reduce was originally coded directly in Lisp 1.5 — but by the end of the 1960s it had been reworked so that it was written in its own “Algol-like” syntax that for the system programmer offered Lisp-like semantics. For example, the following computes Legendre polynomials:
procedure factorial n;
  for i := 1:n product i;
for k := 0:10 do
  write df((x^2-1)^k, x, k)/2^k/factorial k;
@@
Building the system called for a fairly modest size parser in regular Lisp to bootstrap things until the system’s better and more complete version of the same parser expressed in its own language could be installed. This scheme provided significant isolation from dialect variations between competing Lisp implementations — the bootstrap parser could need to be adapted to allow for differences in support for character handling, but then the transformation from textual source code to internal S-expressions readily provided a location where adjustments to the use of Lisp functions could be made. To give a concrete example of the sort of adjustment needed, we note that the Lisp operators map and mapcar that apply a given function to each component of a list took their arguments in differing orders basically depending on whether the Lisp implementation concerned followed East Coast or West Coast traditions. Reduce could adapt for that by including conversion code as part of its parsing phase.
@@
Over time and as the number of computers (and hence Lisp systems) that people wanted to run Reduce on grew this scheme which amounted to source to source translation from the dialect Reduce expected to whatever was locally available became clumsier and more onerous. In response to this Reduce codified a Lisp dialect it would use (calling it “Standard Lisp[18]) so that host systems could arrange to provide a compatibility layer that exposed this agreed version, and the code within Reduce being written based on that could then just run without need for further change. To make this process as easy as possible Standard Lisp mandated only facilities it expected that all plausible Lisps would be able to support, so it did not provide any of the more elaborate additions to Lisp that those who were using the language directly were developing. For instance more or less elaborate loop constructs were present in the Reduce source language but got expanded to simple forms by the time a Lisp “parse tree“ had been generated, so there was no need for the supporting Lisp to provide clever iteration functions.
@@
So the increasing bulk of Reduce being coded in its own language and sitting on a rather minimal base really helped with portability, but two challenges could arise. One was that newer parts of the system sometimes needed facilities not captured in Standard Lisp. Perhaps the best example to quote would be modular arithmetic. This could be provided by merely defining all the necessary functions in terms of ordinary arithmetic with a lot of calls to the remainder function, but really it would be better if implemented at a lower level. Well Reduce could cope with that! It could include portable reference implementations of everything is needed, but a particular implementation could offer built-in and presumably better versions and then tag the relevant function names to cause Reduce to skip instating its reference code. This strategy maintained the policy that the bulk of the code should only very very rarely need to be conditialised based on the platform it was running on.
@@
The second challenge was of course performance. Main-stream Lisps were increasingly focusing their implementation efforts on supporting the full richness of environment expected by users who were coding directly in the parenthesised notation and the optimisations they were delivering were not necessarily the ones most important to Reduce. So as a spin-off activity Reduce first developed its own explicitly portable Lisp compiler, and went on from there is develop of complete Lisp (PSL[14]) explicitly providing (a superset of) Standard Lisp, reasonably easy to move to new machines because it used the portable compiler, and particularly tuned to run Reduce well. PSL was essentially all coded in Lisp as extended with functions that were there to compile into simple operations on machine registers and memory. In broadly the same time-frame Cambridge Lisp[12] was implemented mostly in BCPL but also using a version of the Reduce project’s compiler, and BCPL’s focus on portability meant that ended up being used on a range of the emerging small computers from Amiga, Atari, Acorn and the like. In each case these distilled portability away from the parts of Reduce that supported algebra, and they could both focus on trying to be economical in memory and good in terms of speed for the particular load that Reduce placed on them.
@@
Somewhat later on the use of BCPL became less tenable as C took over the ecological niche it had inhabited and as Unix became a dominant operating system in academia. So Cambridge Lisp was replaced by CSL[40], which was a fully fresh Lisp this time coded in C. A particular feature of this was that its almost entire purpose was to act as a kernel for Reduce rather than as a Lisp for more general use. At the time memory often limited problem solving ability. To save space CSL compiled the Lisp that came out of the Reduce parser into a compact byte-coded form which it could then interpret. One initial insight was that the bytecode interpreter was likely to fit within the caches of the more important computers of the times, and that would lead to the overhead of interpretation being much less severe than one might have at first feared. The second performance trick was to profile a wide range of Reduce applications and on that basis identify the most crucial functions. Those were then compiled from Lisp into equivalent C code that followed the internal CSL conventions and got passed through the system C compiler and linked in as parts of the main CSL executable. By then the leading C compilers (notably gcc) were starting to provide optimisation capabilities distinctly superior to the more generic ones present in the available Lisp compilers. The C code in CSL was kept fairly conservative with a view to keeping it as portable as was reasonably possible.
@@
Because this story spans many years, one should not be surprised to notice that just as BCPL had started to look long in the tooth, at some stage the C code in CSL started to be reworked into C++. The changes there forced further case regarding for instance the difference between immutable strings and updateable arrays of characters, the issues of “strict aliasing”, and the fact that arithmetic overflow on signed integers was deemed undefined. Responding to each of these will have tended to enhance some mix of reliability and portability of the code-base. One of the more concerning issues while that transition was going on was the proper delay time between C++ standards codifying useful new features and the moment that assuming their availability would not seriously compromise attempts to build the code on older machines. While it is often reasonable to expect that personal computers will be kept well up to date with the most recent versions of compilers and other software, major institutional systems can prioritise stability to an extent that to a developer who has not encountered that before it at first seems that they have hardly moved on from the ark. CSL has tried to steer clear of reliance on new language features until they have had plenty of time to propagate to all potential use-sites. This is again in line with the long-standing Reduce philosophy of preferring easy portability over extra and glossy features. Over its lifetime, the storage allocation and garbage collection parts of CSL have been re-written several times, the scheme for preserving a system on disc for later re-loading has been changed also several times and the long integer arithmetic implementation altered rather substantially. But from the outside, all these changes tried to be invisible so that the Reduce system sitting on top did not need to pay any attention, but could benefit from speed improvements or reliability improvements.
@@ 6 Case Study: Aldor
Aldor purposefully retained most of the ideas and syntax of [31] while generalising the type system and providing features to allow clear separation of base language and libraries. Functions and types were first-class values and the language design allowed efficient compilation. Dependent types provided genericity and the structures of modern algebra (such as rings, fields and modules) can be represented naturally as type categories. An optimising compiler for Aldor could generate C code in the Ansi (C90) or K&R C dialects, for stand-alone or linked programs, or Lisp in the Common Lisp, Standard Lisp or Scheme dialects, for embedding in systems based on those. For a fuller description, see [51, 52, 53]. Recall that during its development Axiom had been named Scratchpad II, and similarly Aldor was named $A^{\sharp}$ (and briefly as Axiom-XL).
The main development of Aldor spanned a period of about ten years, from 1984 to 1995. It began as an effort to produce an improved implementation of the language of [31]. It was written in the “Boot” language, described earlier, and it translated algebra code to Lisp to run within the Scratchpad II system.
@@
A wide range of operating systems were supported, including Microsoft DOS and Windows 3.1, IBM OS2 and CMS, Digital Equipment VAX VMS, Apple MAC System 7, and a wide range of Unix variants including AIX PS2, AIX RT, IX RS, AIX 370, AIX ESA (all from IBM), SunOS and Solaris from Sun Microsystems, the Open Software Foundation’s OSF/1, as well as Unicos, Irix, Next, Linux, Ultrix and HP UX. The C environments supported included those from Borland, Metaware, Metroworks, Microsoft, the Free Software Foundation (gcc), four from IBM, DEC’s VMS and Ultrix C, as well as those from Apple, Cray, Hewlett Packard, MIPS and Sun. The architectures supported included Intel 8086 and 80386, Motorola 68000, PowerPC, IBM 370, RT and RS, DEC VAX and AXP, Sun Sparc, MIPS, Cray and Hewlett Packard PA.
@@
The portability layer provided a set of very low-cost abstractions to give a common coding interface for all the supported platforms. This allowed efficient low-level programming while shielding the main body of the code from defects in the various C compilers (of which there were many) and providing common concepts where operating systems differed. This approach necessitated a certain amount of system-specific implementation of low-level functions (for a recent version, 6829 lines of C code for all the operating system variants listed above out of a total of 199,939 lines). This portability layer was the only layer in which platform-specific #ifdefs could occur, and there were about 370 of these.
@@
The portability layer also provided a range of macros to encapsulate code that might have different variants depending on the named properties. For example, ptrDiff(p,q) had different implementations depending on whether or not pointers were canonical. Strictly speaking, the C standard would not allow subtraction of pointers into two different objects, so the ptrDiff operation was needed to support garbage collection. Another example would be that in some (but not all) environments, functions to be used as signal handlers had to have a special modifier in their declaration, and this modifier varied. So a macro SignalModifier would always be used in such declarations, and it would sometimes expand to a keyword and otherwise to nothing. All told, there were about 80 of these macros.
Although the C standard [23] was very new and support for it was lacking on many important platforms, the Aldor source code would be written to the standard. A software tool was written to convert function prototypes and new-style function headers if needed on a given platform, e.g. extern int f(char *s); and int f(char *s) { .. }, to K&R C style, e.g. extern int f(); andint f(s) char *s; { ... }.
@@
Aldor’s run time library provided a conservative garbage collector that made use of the osMemMap function. The root set for garbage collection was obtained by scanning all data regions outside the Aldor heap to find all potential pointers into the Aldor heap, including those from other memory managers.
@@
To give an idea of the amount of portable versus platform-specific code, we note the compiler and run-time system consisted of about 200,000 lines of C code, 1000 lines of 2-level grammar code, and 2,400 lines of Aldor code. Of the C code, about 8000 lines were devoted to the portability layer. About 2000 lines of C code were for software tools were used to build the compiler. These included oldc (to convert Ansi C to K&R C), zacc (to convert 2-level grammar to yacc), and atinlay (to convert the message database to pig latin, for multilingual testing).
Libraries in the Aldor language completed the system, with about 11,000 lines providing basic types and functionality and an additional 41,000 lines providing symbolic mathematical functions, including the $\sum^{IT}$ library [4]. This provided efficient algorithms for sparse and dense linear algebra, polynomial algebra (including factorization, resultants, etc), rational functions and series, as well as difference and differential operators over integers, finite fields and rational numbers.
@@
What we take from this experience is that, with due care and attention to detail, it has been possible to build a highly portable software system for computer algebra that runs across a wide range of diverse platforms by providing a low-level portability layer comprising about 3% of the total code base. The resulting system was able to produce code that ran at speeds comparable to hand-coded C or C++ for the sorts of problems these systems could handle [51, 11]. As we discussed the topics covered here, we were struck by the parallels between the Reduce and Aldor experiences.